# From Quantum Chaos to a Reversed Quantum Disentangled Liquid in a Disorder-Free Spin Ladder

### v1

@@ Abstract
The mechanisms by which isolated interacting quantum systems evade thermalization extend beyond disorder-induced many-body localization, encompassing a growing class of interaction-driven phenomena. We investigate a spin-$\tfrac{1}{2}$ ladder with asymmetric XY leg couplings and tunable Ising interactions on the rungs, and identify the microscopic origin of many-body localization (MBL) in this setting. Through a suite of diagnostics—including entanglement dynamics, fidelity susceptibility, adiabatic gauge potential norms, level-spacing statistics and entropy of eigenstates—we uncover a reentrant progression of dynamical regimes as the rung coupling $J_{z}$ is varied: integrable behavior at $J_{z}=0$, quantum chaos at intermediate $J_{z}$, and a robust nonthermal regime at strong coupling. In the latter regime, we demonstrate the emergence of a reversed quantum disentangled liquid (reversed-QDL), where the light species thermalizes while the heavy species remains localized. The strong-coupling limit further yields emergent local integrals of motion anchored in a fixed-point structure, providing a microscopic origin of the observed quasi-MBL dynamics. These results establish reversed-QDL as a distinct, disorder-free route to nonergodicity and broaden the classification of dynamical phases in quantum matter.
@@ I Introduction
QDLs represent a class of nonthermal states in which subsystems with different dynamical scales decouple entanglement-wise: typically, a “heavy” species thermalizes while a “light” species remains localized. Originally proposed in systems with mass imbalance or internal constraints, QDLs provide a framework for understanding partial ergodicity in constrained quantum dynamics [22, 23, 24, 25]. In parallel, quasi–many-body localization (quasi-MBL)—characterized by slow entanglement growth and long-lived memory retention in the absence of quenched disorder—has been observed in spin ladders and other constrained, disorder-free settings [26, 27, 28, 29, 30, 31]. We stress that by quasi-MBL we refer to disorder-free, finite-size localization-like dynamics, which are distinct from true MBL driven by randomness.
@@ II Model
It would be interesting to mention that the XY Hamiltonian of each leg can be mapped to a spinless free fermion model via a Jordan-Wigner transformation [49, 50]. Let us define $c_{\ell}^{\dagger}$ ($c_{\ell}$) as the creation (annihilation) of a fermion on site $\ell$ on the lower leg and the corresponding ones on the upper leg as $\tilde{c}_{\ell}^{\dagger}$ ($\tilde{c}_{\ell}$). According to the Jordan-Wigner transformation the Pauli operators on the lower leg are defined by
$\tau^{+}_{\ell}$ $=$ $c_{\ell}^{\dagger}e^{(i\pi\sum_{j<\ell}n_{j})},$
$\tau^{-}_{\ell}$ $=$ $e^{(-i\pi\sum_{j<\ell}n_{j})}c_{\ell},$ (2)
$\tau^{z}_{\ell}$ $=$ $2(n_{\ell}-\frac{1}{2})\;\;,\;\;n_{\ell}=c_{\ell}^{\dagger}c_{\ell}.$
An analogous mapping is defined between the Pauli operators $\sigma$ and the fermions $\tilde{c}_{\ell}^{\dagger},\tilde{c}_{\ell}$ (on the upper leg including a Klein factor), where $\{c,\tilde{c}\}=0$. The Hamiltonian of the ladder will be written in terms of fermion operators as
@@
The fermion representation of Hamiltonian (3) shows that the ladder is composed of two free fermion chains coupled via density-density interaction between aligned sites across each rung. It has been mentioned in [26] that for $J^{\prime}=0$, the density of fermions on the upper chain (or equivalently the spin configurations of $\sigma$’s) plays the role of a quenched disorder for the lower chain fermions, leading to (single particle) Anderson localization [51] at strong disorder. Upon taking into account the effect of small $J^{\prime}$ ($J^{\prime}\ll J$), the slow dynamics of heavy fermions ($\tilde{c}$) could act as slow variation of disorder landscape in an interacting fermion system. The long-time dynamics of the model shows a slow decay of polarization [26, 52], which mimics an MBL-like dynamics dubbed quasi-MBL regime. The signature of the latter phase is decoded in the slow (logarithmic) growth of entanglement entropy as shown in Fig. 2 for large Ising coupling ($J_{z}=10,20$). We are going to uncover the mechanism of slow dynamics and investigate the effect of Ising interaction on the dynamical behavior of the model within several diagnostics approaches.
@@ IV.1 Fidelity susceptibility and adiabatic gauge potential norm
$\chi_{\mathrm{typ}}=e^{\zeta}\;;\;\;\zeta\;=\;\langle\log(\chi_{n})\rangle,$ (6)
@@
The inset of Fig. 3 shows the rescaled norm $\|A_{\lambda}\|^{2}/L$ versus $L$ on a logarithmic axis. Fitting $\ln(\|A_{\lambda}\|^{2}/L)=\alpha(J_{z})L+const.$ over L=5,,9L=5,...,9 yields $\alpha=\{0.050,\,2.207,\,0.405,\,0.329,\,0.043\}$ for $J_{z}=\{0,\,10^{-2},\,10^{-1},\,1,\,10\}$, so $\|A_{\lambda}\|^{2}/L\sim e^{\alpha(J_{z})L}$. so that compactly $\|A_{\lambda}\|^{2}/L\sim e^{\alpha(J_{z})L}$.
@@ IV.2 Entropy of eigenstates
We have also computed the mean level spacing as a function of $J_{z}$ (see Appendix A), which demonstrates the crossover from Gaussian statistics at small $J_{z}$ to Poisson-type statistics for large $J_{z}$, consistent with the other diagnostics.
@@ V.1 Measured Entanglement Entropy: QDL Diagnostic
In addition to $S_{L/2}$, we compute the measured entropies as previously defined. The entropy of $\tau$ spins conditioned on measured configurations of $\sigma$ spins, denoted by $S_{A}^{(\tau|\sigma)}$, is shown in blue with square markers. This entropy follows a similar temporal profile to $S_{L/2}$ up to $t_{2}$, and for $t>t_{2}$ displaying logarithmic growth and rapid saturation at $t_{3}$. This behavior suggests that the light species (i.e., $\tau$ spins) thermalizes quickly and does not retain memory of observables, contrary to the expectations of the QDL paradigm (see Appendix B).
@@ V.2 Emergent XXZ Dynamics in the strong coupling limit
The first-order contribution vanishes due to orthogonality with the excited sector, so we proceed with second-order Brillouin–Wigner perturbation theory [60, 61, 62, 63, 64, 65]. The resulting effective Hamiltonian is (for details see Appendix C):
@@
$-(\frac{J^{2}+J^{\prime 2}}{J_{z}})\sum_{l=1}^{L}\hat{\Phi}_{\ell}^{z}\hat{\Phi}_{\ell+1}^{z}+\frac{J^{2}+J^{\prime 2}}{J_{z}}\,\mathbb{1}.$ (20)
@@ Acknowledgements.
Appendix A Level spacing statistics
@@
Appendix B Scaling fits as evidence for a reversed QDL regime
@@
Figure 9: (Color online) (a) Heavy leg ($\sigma$)—measured half-cut entanglement entropy. Main panel show $S^{(\sigma|\tau)}_{L/2}(t)$; the inset reports the terminal–window saturation values $S^{(\sigma|\tau)}_{\rm sat}(L)$ together with the best power-law fit $S_{\rm sat}(L)=a\,L^{\alpha}+b$ with $\alpha\simeq 0.30$, indicating sub-volume growth. (b) Light leg ($\tau$)—measured half-cut entanglement entropy. Main panel show $S^{(\tau|\sigma)}_{L/2}(t)$; the inset shows $S^{(\tau|\sigma)}_{\rm sat}(L)$ with the best linear fit $S_{\rm sat}(L)=aL+b$, consistent with volume-law scaling, where error bars display the one–standard-deviation spread of the terminal window used to define $S_{\rm sat}$. In both panels results are shown for $L=4,6,8$ with periodic boundary conditions (PBC). The model parameters are $J=1$, $J^{\prime}=10^{-3}$, $J_{z}=10$, and all calculations are performed in the symmetry sector $\mathcal{S}^{z}_{\sigma}=\mathcal{S}^{z}_{\tau}=0$ (i.e., $\mathcal{S}^{z}$ per leg =0=0).
@@
From the long–time saturation values [Fig. 9(a)], a two–parameter power law $S^{(\sigma|\tau)}_{\rm sat}(L)=aL^{\alpha}+b$ is decisively preferred by information–theoretic criteria over logarithmic and linear alternatives. Using BIC we find $\mathrm{BIC}_{\rm pow}=-66.45$, $\mathrm{BIC}_{\log}=-25.47$, and $\mathrm{BIC}_{\rm lin}=-20.61$ (with comparable parameter counts), while the fits have $R^{2}_{\rm pow}=1.000$ vs. $R^{2}_{\log}=0.997$ and residual sums of squares $\mathrm{RSS}_{\rm pow}=3.47\times 10^{-10}$ vs. $\mathrm{RSS}_{\log}=2.97\times 10^{-4}$. Leave–one–out CV–MSE is slightly smaller for the logarithmic fit ($4.53\times 10^{-4}$ vs. $7.30\times 10^{-4}$), but the BIC differences in favor of the power law are decisive: $\Delta\mathrm{BIC}_{\rm pow,log}\sim 41$. The best fit yields $\alpha\simeq 0.30$ (robust across tail windows, $\alpha\in[0.22,0.48]$), establishing sub–volume growth of the measured entropy in the heavy sector (see Table 1).
@@
Appendix C Effective Two-Rung Theory in the Strong Coupling Limit

### v2

# Crossover from Quantum Chaos to a Reversed Quantum Disentangled Liquid in a Disorder-Free Spin Ladder

@@ Abstract
The mechanisms by which isolated interacting quantum systems evade thermalization extend beyond disorder-induced many-body localization, encompassing a growing class of interaction-driven phenomena. We investigate a spin-$\tfrac{1}{2}$ ladder with asymmetric XY leg couplings and tunable Ising interactions on the rungs, and identify the microscopic origin of quasi many-body localization (quasi-MBL) in this setting. Through a suite of diagnostics—including entanglement dynamics, fidelity susceptibility, adiabatic gauge potential norms, level-spacing statistics and entropy of eigenstates—we uncover a reentrant progression of dynamical regimes as the rung coupling $J_{z}$ is varied: integrable behavior at $J_{z}=0$, quantum chaos at intermediate $J_{z}$, and a robust nonthermal regime at strong coupling. In the latter regime, we demonstrate the emergence of a reversed quantum disentangled liquid (reversed-QDL), where the light species thermalizes while the heavy species remains localized. The strong-coupling limit further yields emergent local integrals of motion anchored in a fixed-point structure, providing a microscopic origin of the observed quasi-MBL dynamics. These results establish reversed-QDL as a distinct, disorder-free route to nonergodicity and broaden the classification of dynamical phases in quantum matter.
@@ I Introduction
QDLs represent a class of nonthermal states in which subsystems with different dynamical scales decouple entanglement-wise: typically, a “heavy” species thermalizes while a “light” species remains localized. Originally proposed in systems with mass imbalance or internal constraints, QDLs provide a framework for understanding partial ergodicity in constrained quantum dynamics [22, 23, 24, 25]. In parallel, quasi–many-body localization (quasi-MBL)—characterized by slow entanglement growth and long-lived memory retention in the absence of quenched disorder—has been observed in spin ladders and other constrained, disorder-free settings [26, 27, 28, 29, 30, 31, 32]. We stress that by quasi-MBL we refer to disorder-free, finite-size localization-like dynamics, which are distinct from true MBL driven by randomness.
@@ II Model
It would be interesting to mention that the XY Hamiltonian of each leg can be mapped to a spinless free fermion model via a Jordan-Wigner transformation [50, 51]. Let us define $c_{\ell}^{\dagger}$ ($c_{\ell}$) as the creation (annihilation) of a fermion on site $\ell$ on the lower leg and the corresponding ones on the upper leg as $\tilde{c}_{\ell}^{\dagger}$ ($\tilde{c}_{\ell}$), where $\{c,\tilde{c}\}=0$. The details of transformation is given in Appendix A. The Hamiltonian of the ladder will be written in terms of fermion operators as
@@
The fermion representation of Hamiltonian (2) shows that the ladder is composed of two free fermion chains coupled via density-density interaction between aligned sites across each rung. It has been mentioned in [26] that for $J^{\prime}=0$, the density of fermions on the upper chain (or equivalently the spin configurations of $\sigma$’s) plays the role of a quenched disorder for the lower chain fermions, leading to (single particle) Anderson localization [52] at strong disorder. Upon taking into account the effect of small $J^{\prime}$ ($J^{\prime}\ll J$), the slow dynamics of heavy fermions ($\tilde{c}$) could act as slow variation of disorder landscape in an interacting fermion system. The long-time dynamics of the model shows a slow decay of polarization [26, 53], which mimics an MBL-like dynamics dubbed quasi-MBL regime. The signature of the latter phase is decoded in the slow (logarithmic) growth of entanglement entropy as shown in Fig. 2 for large Ising coupling ($J_{z}=10,20$). We would like to stress that all numerical simulations in this article are conducted on the spin Hamiltonian defined in Eq. (1). We are going to uncover the mechanism of slow dynamics and investigate the effect of Ising interaction on the dynamical behavior of the model within several diagnostics approaches.
@@ IV.1 Fidelity susceptibility and adiabatic gauge potential norm
$\chi_{\mathrm{typ}}=e^{\zeta}\;;\;\;\zeta\;=\;\langle\ln(\chi_{n})\rangle,$ (5)
@@
The inset of Fig. 3 shows the rescaled norm $\|A_{\lambda}\|^{2}/L$ versus $L$ on a logarithmic axis. Fitting $\ln(\|A_{\lambda}\|^{2}/L)=\alpha(J_{z})L+const.$ over L=5,,9L=5,...,9 yields $\alpha=\{0.050,\,2.207,\,0.405,\,0.329,\,0.043\}$ for $J_{z}=\{0,\,10^{-2},\,10^{-1},\,1,\,10\}$, so that compactly $\|A_{\lambda}\|^{2}/L\sim e^{\alpha(J_{z})L}$.
@@ IV.2 Entropy of eigenstates
We have also computed the mean level spacing as a function of $J_{z}$ (see Appendix B), which demonstrates the crossover from Gaussian statistics at small $J_{z}$ to Poisson-type statistics for large $J_{z}$, consistent with the other diagnostics.
@@ V.1 Measured Entanglement Entropy: QDL Diagnostic
In addition to $S_{L/2}$, we compute the measured entropies as previously defined. The entropy of $\tau$ spins conditioned on measured configurations of $\sigma$ spins, denoted by $S_{A}^{(\tau|\sigma)}$, is shown in blue with square markers. This entropy follows a similar temporal profile to $S_{L/2}$ up to $t_{2}$, and for $t>t_{2}$ displaying logarithmic growth and rapid saturation at $t_{3}$. This behavior suggests that the light species (i.e., $\tau$ spins) thermalizes quickly and does not retain memory of observables, contrary to the expectations of the QDL paradigm (see Appendix C).
@@ V.2 Emergent XXZ Dynamics in the strong coupling limit
The first-order contribution vanishes due to orthogonality with the excited sector, so we proceed with the second-order Brillouin–Wigner perturbation theory [61, 62, 63, 64, 65, 66]. The resulting effective Hamiltonian is (for details see Appendix D):
@@
$-(\frac{J^{2}+J^{\prime 2}}{J_{z}})\sum_{l=1}^{L}\hat{\Phi}_{\ell}^{z}\hat{\Phi}_{\ell+1}^{z}+\left(\frac{J^{2}+J^{\prime 2}}{J_{z}}\right)\mathbb{1}.$ (19)
@@ Acknowledgements.
Appendix A Jordan-Wigner transformation
Although we do not numerically simulate the fermionic model of Eq. (2), we briefly explain the Jordan-Wigner transformation of the original spin model (Eq. (1)) to its fermionic counterpart. According to the Jordan-Wigner transformation [50, 51], the Pauli operators on the lower leg ($\tau$-spins) are defined by
$\tau^{+}_{\ell}$ $=$ $c_{\ell}^{\dagger}\;exp(i\pi\sum_{j<\ell}n_{j}),$
$\tau^{-}_{\ell}$ $=$ $exp(-i\pi\sum_{j<\ell}n_{j})\;c_{\ell},$ (20)
$\tau^{z}_{\ell}$ $=$ $2(n_{\ell}-\frac{1}{2})\;\;,\;\;n_{\ell}=c_{\ell}^{\dagger}c_{\ell}.$
Accordingly, we get
$\tau^{+}_{\ell}\tau^{-}_{\ell+1}+\tau^{-}_{\ell}\tau^{+}_{\ell+1}=c_{\ell}^{\dagger}c_{\ell+1}+c_{\ell+1}^{\dagger}c_{\ell},$ (21)
for $\ell=1,\cdots,L-1$, except the boundary terms. The corresponding boundary terms for the lower leg are given by:
$\tau^{+}_{L}\tau^{-}_{1}+\tau^{-}_{L}\tau^{+}_{1}=-F(c_{L}^{\dagger}c_{1}+c_{1}^{\dagger}c_{L}),$ (22)
where $F$ is the Klein factor
$F=exp(i\pi\hat{N})\;\;\;;\;\;\;\hat{N}=\sum_{\ell=1}^{L}n_{\ell}.$ (23)
A similar transformation for the upper leg ($\sigma$-spins) is given by the following equations, where the inclusion of the Klein factor is responsible for preserving the commutation relations of the original spin operators,
$\sigma^{+}_{\ell}$ $=$ $\tilde{c}_{\ell}^{\dagger}\;exp(i\pi\sum_{j<\ell}\tilde{n}_{j})\;F,$
$\sigma^{-}_{\ell}$ $=$ $F^{\dagger}\;exp(-i\pi\sum_{j<\ell}\tilde{n}_{j})\;\tilde{c}_{\ell}$, (24)
$\sigma^{z}_{\ell}$ $=$ $2(\tilde{n}_{\ell}-\frac{1}{2})\;\;,\;\;\tilde{n}_{\ell}=\tilde{c}_{\ell}^{\dagger}\tilde{c}_{\ell}$.
After some calculations, we arrive at the following boundary term for the upper leg:
$\sigma^{+}_{L}\sigma^{-}_{1}+\sigma^{-}_{L}\sigma^{+}_{1}=-\tilde{F}(\tilde{c}_{L}^{\dagger}\tilde{c}_{1}+\tilde{c}_{1}^{\dagger}\tilde{c}_{L}),$ (25)
where,
$\tilde{F}=exp(i\pi\hat{\tilde{N}})\;\;\;;\;\;\;\hat{\tilde{N}}=\sum_{\ell=1}^{L}\tilde{n}_{\ell}$. (26)
Both $F$ and $\tilde{F}$ yield $\pm 1$ depending on the parity of fermion occupations of the lower and upper legs, respectively. The boundary terms in Eqs.(22) and (25) demonstrate that an odd (even) parity of fermion occupations in either the legs of ladder necessitates a periodic (anti-periodic) boundary term for the fermionic model. Finally, we arrive at the fermionic Hamiltonian, Eq. (2).
Appendix B Level spacing statistics
@@
Appendix C Scaling fits as evidence for a reversed QDL regime
@@
[…]
Figure 9: (Color online) (a) Heavy leg ($\sigma$)—measured half-cut entanglement entropy. Main panel shows $S^{(\sigma|\tau)}_{L/2}(t)$; the inset reports the terminal-window saturation values $S^{(\sigma|\tau)}_{\rm sat}(L)$ together with the best power-law fit $S_{\rm sat}(L)=aL^{\alpha}+b$ with $\alpha\simeq 0.30$, indicating sub-volume growth. (b) Light leg ($\tau$)—measured half-cut entanglement entropy. Main panel shows $S^{(\tau|\sigma)}_{L/2}(t)$; the inset displays $S^{(\tau|\sigma)}_{\rm sat}(L)$ with the best linear fit $S_{\rm sat}(L)=aL+b$, consistent with volume-law scaling. Error bars denote the one–standard-deviation spread of the terminal window used to define $S_{\rm sat}$. (c) Long-time saturation values of the measured half-cut entanglement entropy versus $J_{z}$ for $L=8$, showing the separation between $S^{(\tau|\sigma)}_{L/2}$ and $S^{(\sigma|\tau)}_{L/2}$ at strong coupling. Results in (a) and (b) are shown for $L=4,6,8$ with periodic boundary conditions. Model parameters are $J=1$, $J^{\prime}=10^{-3}$, $J_{z}=10$, and calculations are performed in the symmetry sector $\mathcal{S}^{z}_{\sigma}=\mathcal{S}^{z}_{\tau}=0$.
@@
From the long–time saturation values [Fig. 9(a)], a two–parameter power law $S^{(\sigma|\tau)}_{\rm sat}(L)=aL^{\alpha}+b$ is decisively preferred by information–theoretic criteria over logarithmic and linear alternatives. Using BIC we find $\mathrm{BIC}_{\rm pow}=-66.45$, $\mathrm{BIC}_{\ln}=-25.47$, and $\mathrm{BIC}_{\rm lin}=-20.61$ (with comparable parameter counts), while the fits have $R^{2}_{\rm pow}=1.000$ vs. $R^{2}_{\ln}=0.997$ and residual sums of squares $\mathrm{RSS}_{\rm pow}=3.47\times 10^{-10}$ vs. $\mathrm{RSS}_{\ln}=2.97\times 10^{-4}$. Leave–one–out CV–MSE is slightly smaller for the logarithmic fit ($4.53\times 10^{-4}$ vs. $7.30\times 10^{-4}$), but the BIC differences in favor of the power law are decisive: $\Delta\mathrm{BIC}_{\rm pow,log}\sim 41$. The best fit yields $\alpha\simeq 0.30$ (robust across tail windows, $\alpha\in[0.22,0.48]$), establishing sub–volume growth of the measured entropy in the heavy sector (see Table 1).
@@
For completeness, we examine the dependence of the long-time saturation values of the measured half-cut entanglement entropy on the inter-leg coupling $J_{z}$ at fixed system size $L=8$. The saturation values are extracted using the terminal window averaging procedure defined in Eq. (28) over the interval $t\in[10^{11},10^{12}]$. As shown in Fig. 9(c), both measurement protocols yield comparable entropies for weak coupling ($J_{z}\ll 1$), close to the thermal (Page) value expected for a half bipartition of this system, $S_{\mathrm{Page}}\simeq\ln(2^{L/2})-\tfrac{1}{2}\approx 2.27$ for $L=8$. However, as $J_{z}$ increases beyond $1$, a distinct separation emerges: $S^{(\tau|\sigma)}_{L/2}$ remains relatively larger, while $S^{(\sigma|\tau)}_{L/2}$ is strongly suppressed. This behavior is consistent with the finite-size scaling analysis presented above, indicating that the light ($\tau$) sector remains thermal with volume-law measured entanglement, while the heavy ($\sigma$) sector exhibits strongly reduced, sub-volume entanglement consistent with partial localization, as expected in the reversed quantum disentangled liquid regime.
@@
Appendix D Effective Two-Rung Theory in the Strong Coupling Limit